OpenStudy (anonymous):

Simplify. Assume that no variable equals 0. (5x^3y^-5)(4xy^3)

OpenStudy (anonymous):

Remember that when multiplying variables with exponents, when you multiply with the same variable, say x^2*x^5, you simply add the exponents, so x^2*x^5 = x^7. When they are divided, you do the same thing except you SUBTRACT the exponents. Also, remember that when you have a negative exponent, for example x^-5, the negative sign means that the variable is flipped: so, x^-5 = 1/x^5